Jim wants to buy cement to anchor two posts for a swing he is building. He has dug 2 round holes that 12 inches in diameter and 40 inches deep. He wants to fill each with concrete around a rectangular supporting beam that is 3.5 inches by 5.5 inches. Each 80-pound bag of concrete mixes with water to make 0.6 cubic fee of concrete. How many bags of concrete must he buy for the job?

if you would explain how you got the answer to this question that would be awesome. thanks!

Found 2 solutions by stanbon, rajagopalan:
Answer by stanbon(48522)

Jim wants to buy cement to anchor two posts for a swing he is building. He has dug 2 round holes that 12 inches in diameter and 40 inches deep. He wants to fill each with concrete around a rectangular supporting beam that is 3.5 inches by 5.5 inches. Each 80-pound bag of concrete mixes with water to make 0.6 cubic fee of concrete. How many bags of concrete must he buy for the job?
---------------------
Volume of the post: 3.5*5.5*40 = 770 cu. in.
Volume of the holes: 2*(pi)6^2*40 = 9047.786842 cu. in
-----------------------
Volume he needs to fill with concrete:
9047.786842-770 = 8277.786842 cu. in
-----------------------
1 cu. ft of of concrete = 12*12*12 = 1728 cu. in.
0.6 cu. ft. = 0.6*1728 = 1036.8 cu. in.
-----------------------------
# of bags = 8277.786842/1036.8 = 7.98 bags
-----------------------------
He needs to buy 8 bags.
=============
Cheers,
Stan H.

Answer by rajagopalan(148)

volume of 2 pits =2xpixrxrxh=2x(22/7)x(12/2)x(12/2)x40= 9051.42 cu in
volume of pole=2(LxBxH)=2x 770=1540 cu in
volume of cement needed = vol of pits-vol of poles=9051-1540=8281.42cu inch
volume of cement needed = 8281.42cu inch/(12x12x12) cu ft = 4.34 cu ft
no of bags = Vol of cement needed /one bag capacity= 4.34/0.6=7.2
Say 8 bags ( the next integer)
*
He has to buy 8 bags of cement
okay